OpenStudy (anonymous):

1. The cost, in millions of dollars, to remove x % of pollution in a lake modeled by a. What is the cost to remove 75% of the pollutant? b. What is the cost to remove 90% of the pollutant? c. What is the cost to remove 99% of the pollutant? d. For what value is this equation undefined?

OpenStudy (anonymous):

modeled by c= 6000/200-2x

OpenStudy (anonymous):

start replacing the variable \(x\) by the numbers 75, 90, 99

OpenStudy (anonymous):

for example, the first one is \[C=\frac{6000}{200-2\times 75}=\frac{6000}{200-150}=\frac{6000}{50}=120\]

OpenStudy (anonymous):

as for part d it is undefined when \(x=100\) because \(200-2\times 100=0\) and you cannot divide by zero